OpenStudy (bloomlocke367):

A logarithmic function is the inverse of an exponential function. Can someone explain this to me?

OpenStudy (bloomlocke367):

@Nnesha

OpenStudy (bloomlocke367):

@bibby

OpenStudy (irishboy123):

(1) p = ln q (2) e^p = e^(ln(q)) = q (3) [AND: just (1) in reverse] ln q = p see this: http://www.mathsisfun.com/sets/function-inverse.html

OpenStudy (bibby):

y=e^x->x=e^y solve for y

OpenStudy (bloomlocke367):

I forgot how to do that... o.o

OpenStudy (triciaal):

the log of a number is the power to which the base must be raised to get that number

Nnesha (nnesha):

to cancel out "e "you have to take ln both side for example \[\large\rm ln e^x = x\] \[\ln 1 =0\] \[\ln e^7 =7\]

OpenStudy (bloomlocke367):

what's ln?

Nnesha (nnesha):

natural logarithm ;)

Nnesha (nnesha):

same rules for ln and log

OpenStudy (bloomlocke367):

so what's that? is it different from just log?

OpenStudy (irishboy123):

@bibby \[y = e ^ {x} => x = e ^{y} => y = e ^ {e ^{ y}} \]

OpenStudy (bloomlocke367):

Okay. This is confusing me

Nnesha (nnesha):

nope same properties for ln but you have to use ln when you are dealing with "e" otherwise \[\rm ln x + \ln y= \ln(x \times y)~~~~ logx + \log y = \log (x \times y)\]

OpenStudy (bloomlocke367):

why?

Nnesha (nnesha):

why what ?? why ln cancel out e ??

OpenStudy (bloomlocke367):

why do you have to use ln with e?

OpenStudy (triciaal):